A student dissolved 0.44g of copper wire and followed all steps described in the copper cycle experiment to obtain solid copper. After performing the series of reactions a student weighed a total mass of 24.61g on the scale. This mass included the filter paper mass of 0.23g and the watch glass of 24.19g. Calculate the % recovery for this experiment. %recovery= final mass of copper grams divided by initial mass of copper grams
Answer #1

Initial mass of Copper = 0.44 g

Mass of filter paper = 0.23 g

Mass of watch glass = 24.19 g

Mass of filter paper and watch glass = 0.23 + 24.19 = 24.42 g

Total mass = 24.61 g

Mass of copper recovered = 24.61 – 24.42 = 0.19 g

% recovery = Mass of copper recovered /Initial mass of Copper *100

% recovery = 0.19 g / 0.44 g*100 = 43.18182 % = 43.18 %